What is the magnitude of the force on the proton i
Answer #1

a)
F = q*E
=> F = 1.6e-19*e6

=> F = 1.6e-13 N

b)
F = q*v*B*sin(theta)

=> 1.6e-13 = 1.6e-19*1.5e7*0.10*sin(theta)
=> sin(theta) = (1.6e-13)/( 1.6e-19*1.5e7*0.10)
=> sin(theta) = 0.67
=> theta = 42 degrees